OpenStudy (anonymous):

write an equation of a line in point slope form that has a slope of -3, and passes through the point (3,-4)

OpenStudy (blurbendy):

use the formula: y - y1 = slope(x - x1)

OpenStudy (anonymous):

but, i only have one set of points.

OpenStudy (blurbendy):

exactly, so put them in for y1 and x1 accordingly

OpenStudy (anonymous):

ok. so then, y-(-4)=-3(x-3)?

OpenStudy (blurbendy):

yup that simplifies to y + 4 = -3(x - 3)

OpenStudy (blurbendy):

Distribute the -3 on the right side

OpenStudy (anonymous):

y+4=3x+9

OpenStudy (blurbendy):

-3x

OpenStudy (blurbendy):

Now, get y b itself on the left side

OpenStudy (anonymous):

y=-3x+5

OpenStudy (blurbendy):

yup, good job!

OpenStudy (anonymous):

ok, thats in slope intercept form isnt it? not point slope form

OpenStudy (blurbendy):

yes, slope-intercept form is y = mx + b which is what we have

OpenStudy (anonymous):

right. but it says write the equation in point slope form, not slope intercept form.

OpenStudy (anonymous):

blurbendy, i still need your help

OpenStudy (blurbendy):

point slope form is y - y1 = x(x - x1)